Suppose government purchases increase to 1000, what is the new equilibrium income and consumption? (4 Marks)


v. What is the new fiscal deficit/surplus after government spending increases to 1000?



1
Expert's answer
2022-03-20T19:08:41-0400

Solution.

Y = C + I + G

Y = C+ I + 1000


Fiscal deficit = Total expenditure - Total receipts (excluding borrowings).

Fiscal deficit = (Revenue expenditure + Capital expenditure) - (Revenue receipts + Capital receipts excluding borrowings).

Let Y be total receipts

Therefore fiscal deficit = 1000 - Y